OpenStudy (anonymous):

Find the maximum or minimum value of the function. g(x) = 2x(x − 4) + 7

OpenStudy (lucaz):

this is a quadratic function, or parabola, expanding 2x^2-8x+7 from the quadratic formula you get the vertex as -b/2a, here a is 2(from 2x^2) and b is -2(from -8x), substitute them: -(-8)/2(2)=8/4=2, vertex is (2,y), substituting x by 2 you get 2(2)(2-4)+7=-1, vertex=(2,-1) since this parabola has a positive leading coefficient, +2x^2, it opens upward, so the vertex shows the minimum value

OpenStudy (lucaz):

lol, b is -8, not -2

OpenStudy (anonymous):

oh ok thanks!11

OpenStudy (lucaz):

you're welcome